I really don't understand these problems and need help.
State the location of the foci for this hyperbola:
(y-3)^2/11 - (x+1)^2/5 =1

Answer by venugopalramana(3286)

I really don't understand these problems and need help.
State the location of the foci for this hyperbola:
(y-3)^2/11 - (x+1)^2/5 =1
COMPARE WITH STD.EQN.
STD.EQN.IS
(Y-K)^2/B^2 - (X-H)^2/A^2 =1
ITS CENTRE IS (H,K)......HERE....(-1,3)....H=-1,K=3
LENGTH OF CONJUGATE AXIS = 2A ..........HERE....A=SQRT(5)
LENGTH OF TRANSVERSE AXIS = 2B...........HERE....B=SQRT(11)
ECCENTRICITY = E = SQRT[(A^2+B^2)/B^2]...HERE....SQRT[(5+11)/11]=4/SQRT(11)
FOCI ARE (H,K+ OR - B*E)
[-1,3+SQRT(11)*4/SQRT(11)]=(-1,7) AND
[-1,3-SQRT(11)*4/SQRT(11)]=(-1,-1)